Two resistances 4 Ω and 8 Ω are connected in series. A voltage of 24 volts is applied across the combination. Find:

a. The total resistance, and

b. The voltage across each resistor



1
Expert's answer
2021-05-31T02:08:09-0400

Total resistance = 1/4 + 1/8

= 2+1/8 = 3/8

= 8/3

= 2.667 ohms

voltage = 24/2.667

= 8.9999V